I don't understand how to solve this problem. Thank you.
It says solve and check the quadratic equation below by taking the square root of both sides. Express irrational answers in rational form.
(x-10)^2=4
thank you so much

Answer by Theo(13342)

equation is:
(x-10)^2=4
take the square root of both sides of the equation to get:
x-10 = sqrt(4) which becomes:
x-10 = 2
add 10 to both sides of the equation to get:
x = 12
substitute for x in the original equation to get:
(12-10)^2 = 4
simplify to get:
2^2 = 4
simplify further to get:
4 = 4
x = 12 is confirmed as good.
that's your answer.